A large fireplace in a hotel has a rectangular opening with a height of 5 feet and a diagonal of 13 feet, What is the width, X, of the fireplace opening?

Answer by nerdybill(6963)

A large fireplace in a hotel has a rectangular opening with a height of 5 feet and a diagonal of 13 feet, What is the width, X, of the fireplace opening?
.
We can apply Pythagorean Theorem:
Let x = width
then
x^2 + 5^2 = 13^2
x^2 + 25 = 169
x^2 = 144
x = 12 feet